# EEG Emotion Copilot: Optimizing Lightweight LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

### v2

@@ Abstract
In the fields of affective computing (AC) and brain-machine interface (BMI), the analysis of physiological and behavioral signals to discern individual emotional states has emerged as a critical research frontier. While deep learning-based approaches have made notable strides in EEG emotion recognition, particularly in feature extraction and pattern recognition, significant challenges persist in achieving end-to-end emotion computation, including real-time processing, individual adaptation, and seamless user interaction. This paper presents the EEG Emotion Copilot, a system optimizing a lightweight large language model (LLM) with 0.5B parameters operating in a local setting, which first recognizes emotional states directly from EEG signals, subsequently generates personalized diagnostic and treatment suggestions, and finally supports the automation of assisted electronic medical records. Specifically, we demonstrate the critical techniques in the novel data structure of prompt, model pruning and fine-tuning training, and deployment strategies aiming at improving real-time performance and computational efficiency. Extensive experiments show that our optimized lightweight LLM-based copilot achieves an enhanced intuitive interface for participant interaction, superior accuracy of emotion recognition and assisted electronic medical records generation, in comparison to such models with similar scale parameters or large-scale parameters such as 1.5B, 1.8B, 3B and 7B. In summary, through these efforts, the proposed copilot is expected to advance the application of AC in the medical domain, offering innovative solution to mental health monitoring. The codes will be released at https://github.com/NZWANG/EEG_Emotion_Copilot.
Index Terms:
Lightweight LLM, Model pruning, Model fine-tuning, Emotion recognition, Assisted electronic medical record
I Introduction
@@
THE application of affective computing (AC) in brain-machine interfaces (BMI) is emerging as a key research frontier. Affective computing [1] seeks to identify emotional states by analyzing physiological and behavioral signals, such as electroencephalograms (EEG) [2] [3], heart rate [4], facial expressions [5], voice [6], etc. Advances in this field have expanded the possibilities for BMI technology, particularly in enhancing human-machine interaction [7], [8], and increasing its potential for practical applications in rehabilitation [9], [10] and other areas [11].
@@
The powerful capabilities of deep neural networks (DNNs) have driven rapid advancements in AC for emotion recognition via EEG signals [12], [13], with numerous deep learning approaches proposed [14], [15]. Recently, large language models (LLMs) [16], [17] built on Transformer architecture [18] have demonstrated strong contextual understanding [19]. Models such as EEG-GPT [20] and ChatGPT-BCI [21] focus primarily on brain state identification but have yet to effectively address diverse, human-centered tasks. The integration of LLMs with EEG signals to generate personalized diagnoses, treatment plans, and assisted electronic medical records remains an urgent and complex challenge with the following limitations.
@@
Secondly, EEG signals can be viewed as a form of long text data [29], yet the signal sequences from a single channel often exhibit significant redundancy for machine learning approaches [30], [31] [32], [33]. LLMs require a considerable number of tokens [18] to process this redundancy, resulting in unnecessary computational overhead. Additionally, many AC tasks rely on EEG data from 32, 64, or more channels, further diminishing computational efficiency and presenting challenges for real-time performance. Therefore, effective data compression is critical to optimizing lightweight LLMs for real-time applications.
@@
II RELATED WORK
II-A Lengthy EEG Signal Embedded in Prompt
For specific tasks in LLM prompt engineering, tailored prompts are essential to obtain reasonable answers [34]. Although datasets like Wikitext [25] and Common Crawl [26] contain extensive knowledge, constructing specific datasets is necessary for optimal performance in particular tasks. In most studies on AC, researchers focus on distinguishing emotions using convolutional neural networks (CNNs) or recurrent neural networks (RNNs) [35], with EEG signals primarily used to assess model quality [36], [37]. Recently, LLMs should split EEG sequences into tokens for processing, with computational and memory requirements increasing exponentially with input length [38]. This poses challenges for LLMs deployment, especially on low-resource devices or in real-time scenarios. Although data chunking strategies [39] can split long sequences for training and deployment, the EEG data length significantly impacts reasoning effectiveness [40], [41]. Therefore, a reasonable data compression method is necessary to LLMs.
II-B Model Pruning in Deep Learning
@@
II-C Assisted Electronic Medical Records
Electronic medical records (EMRs) date back to the 1960s [56]. With advancements in computer science and artificial intelligence (AI), EMRs have evolved from merely storing and managing patient information to incorporating more complex functionalities. Currently AI models are increasingly integrated into diagnosing mental illnesses and detecting internal organ pathologies, enhancing the capabilities of assisted EMRs. The sample of ”assisted electronic medical record” is shown in Fig. 2, where the intelligent copilot can process the multimodal diagnostic data and produce accurate results. In this study, we developed a new EEG Emotion Copilot that leverages a lightweight local LLM to perform emotion recognition based on EEG signals, and to further generate the corresponding assisted EMR.
@@ II-C Assisted Electronic Medical Records
III METHOD
III-A Data Structures of Prompt in EEG Emotion Copilot
Fine-tuned LLMs often achieve strong performance to generate end-to-end outputs, particularly for simple classification tasks. For instance, distinguishing whether a subject’s emotional state is positive or negative based on EEG data is relatively straightforward. However, distinguishing more nuanced emotions, such as nine-class emotions in FACED dataset [28], or accurately identifying emotions in real-world scenarios, remains challenging.
For Wikitext or similar datasets, these datasets often consist of declarative sentences (left part of Fig. 3), but for tasks like logical reasoning, data are stored as question-answer pairs, as shown in the middle part of Fig. 3. Despite these advancements, current data structures encounter significant limitations when applied to EEG-based emotion diagnosis and treatment planning. These challenges arise particularly in integrating multimodal data and addressing task-specific requirements. To address these gaps, we propose a novel structure leveraging a question-answer system tailored to EEG-related tasks.
@@ III-A Data Structures of Prompt in EEG Emotion Copilot
In the data structure of the proposed EEG Emotion Copilot, we initialize the system with the prompt showed in the right part of Fig. 3: ”You are an EEG emotion analyzer. I will input the patient’s personal information and EEG signals from specific electrode positions. Please infer the patient’s current emotional state and provide a detailed diagnosis along with personalized treatments based on your knowledge base.” Following initialization, we input the subject’s demographic data, including gender, age, and facial features, to assess their emotional state. Next, the acquired EEG signals are processed. As illustrated in Fig. 1, preprocessing steps such as artifact removal and re-referencing are applied to the signals, followed by wavelet transformation for signal compression. These processed signals are then used as input. Finally, the structure is completed by adding emotion labels and corresponding treatment plans in the second and third sections, respectively.
III-B Model Pruning in EEG Emotion Copilot
@@
$\mathcal{L}=Num(Qwen2\_layer)=Num(Embedding\_layer)$ (1)
$+Num(Qwen2Decoder\_layer)+Num(Lm\_head)=25.$
The input and output of each layer are represented as $f_{i}^{-}$ and $f_{j}^{+}$, respectively, where $0\leq i\leq 25,0\leq j\leq 25$ and $i,j\in\mathbb{N}^{+}$.
Next, we construct a layer dependency graph $\mathcal{F}$ according to [57]. Based on the symmetric mechanism of input $f_{i}^{-}$ and output $f_{j}^{+}$, the dependency graph $\mathcal{F}$ is defined as a $\mathcal{L}\times 2\mathcal{L}$ symmetric matrix, specifically as follows:
$\mathcal{F}\left(f_{i}^{-},f_{j}^{+}\right)$ $=\mathcal{F}\left(f_{j}^{+},f_{i}^{-}\right)$ (2)
$=$ $\begin{cases}1,&\text{if }f_{i}^{-}\leftrightarrow f_{j}^{+}\text{ or }\\ &\;\left(i=j\text{ and }\operatorname{sch}\left(f_{i}^{-}\right)=\operatorname{sch}\left(f_{j}^{+}\right)\right),\\ 0,&\text{otherwise}\end{cases}$
where $\text{sch}(p)$ operation is used to determine which parameters in the parameter group $p$ need to be retained. Formula 2 returns a boolean value, where $\text{sch}(f_{i}^{+})=\text{sch}(f_{j}^{-})$ indicates that the two parameter groups are subject to the same pruning scheme.
@@
$\small\mathcal{I}(\text{embed\_tokens})=\sum_{w\in\text{embed\_tokens}}\|w\|_{2}^{2},$ (3)
$\footnotesize\mathcal{I}(s_{i})=\sum_{w\in s_{i}}\|w\|_{2}^{2},\;S=\{s_{i}\}=\{q_{\text{proj}},k_{\text{proj}},v_{\text{proj}},o_{\text{proj}}\},$ (4)
$\footnotesize\mathcal{I}(m_{i})=\sum_{w\in m_{i}}\|w\|_{2}^{2},\;M=\{m_{i}\}=\{gate_{\text{proj}},up_{\text{proj}},down_{\text{proj}}\},$ (5)
$\small\mathcal{I}(\text{lm\_head})=\sum_{w\in\text{lm\_head}}\|w\|_{2}^{2}.$ (6)
@@
1:Input: Dependency graph $\mathcal{F}$, parameter group set $\mathcal{P}$
2:Output: Group set $\mathcal{P}^{\prime}$
3:Initialize $\mathcal{P}^{\prime}\leftarrow\emptyset$
4:for all $i$in $\mathcal{L}$  do
5:     Create new group $p\leftarrow\{k\}$
6:     Expand group $p$:
7:     while there exists an unvisited node $j$ such that $\mathcal{F}$$(f_{i}^{-},f_{j}^{+})=1$ do
8:         $p\leftarrow p\cup\{j\}$
9:         Mark $j$ as visited
10:     end while
11:     $\mathcal{P}^{\prime}\leftarrow\mathcal{P}^{\prime}\cup p$
12:end for
13:return $\mathcal{P}^{\prime}$
III-C Model Fine-tuning in EEG Emotion Copilot
@@
Specifically, the model’s fine-tuning process on the specific EEG dataset is illustrated in Fig. 6, dividing the training process into five steps, each with three epochs. In steps 1–4, we used 250 Hz data from 55 subjects of FACED [28]. Interestingly, the model’s performance in step 3 surpassed that in step 4, despite minimal differences in training_loss, suggesting that excessive training may be unnecessary and even detrimental for lightweight models. In step 5, 1000 Hz data from 68 FACED subjects pre-trained on step 4 yielded richer responses, confirming the model’s capability to learn the data structure and generate task-specific answers effectively.
III-D Model Deployment in EEG Emotion Copilot
Given the various model deployment tools, such as llama.cpp, which can quantify the model based on the specific scenario and convert it into gguf format, it can be directly utilized through ollama, llmstudio, gradio or other tools. This process with gradio ultimately results in the model deployment part illustrated in Fig. 1.

### v3

@@ Abstract
In the fields of affective computing (AC) and brain-computer interface (BCI), the analysis of physiological and behavioral signals to discern individual emotional states has emerged as a critical research frontier. While deep learning-based approaches have made notable strides in EEG emotion recognition, particularly in feature extraction and pattern recognition, significant challenges persist in achieving end-to-end emotion computation, including rapid processing, individual adaptation, and seamless user interaction. This paper presents the EEG Emotion Copilot, a system optimizing a lightweight large language model (LLM) with 0.5B parameters operating in a local setting, which first recognizes emotional states directly from EEG signals, subsequently generates personalized diagnostic and treatment suggestions, and finally supports the automation of assisted electronic medical records. Specifically, we demonstrate the critical techniques in the novel data structure of prompt, model pruning and fine-tuning training, and deployment strategies aiming at improving performance and computational efficiency. Extensive experiments show that our optimized lightweight LLM-based copilot achieves an enhanced intuitive interface for participant interaction, superior accuracy of emotion recognition and assisted electronic medical records generation, in comparison to such models with similar scale parameters or large-scale parameters such as 1.5B, 1.8B, 3B and 7B. In summary, through these efforts, the proposed copilot is expected to advance the application of AC in the medical domain, offering innovative solution to mental health monitoring. The codes will be released at https://github.com/NZWANG/EEG_Emotion_Copilot.
keywords:
[…]
1 Introduction
@@
The application of affective computing (AC) in brain-computer interfaces (BCI) is emerging as a key research frontier. Affective computing [1] seeks to identify emotional states by analyzing physiological and behavioral signals, such as electroencephalograms (EEG) [2] [3], heart rate [4], facial expressions [5], voice [6], etc. Advances in this field have expanded the possibilities for BCI technology, particularly in enhancing human-machine interaction [7], [8], and increasing its potential for practical applications in rehabilitation [9], [10] and other areas [11].
@@
The powerful capabilities of deep neural networks (DNNs) have driven rapid advancements in AC for emotion recognition via EEG signals [12], [13], with numerous deep learning approaches proposed [14], [15], [16]. Recently, large language models (LLMs) [17], [18] built on Transformer architecture [19] have demonstrated strong contextual understanding [20]. Models such as EEG-GPT [21] and ChatGPT-BCI [22] focus primarily on brain state identification but have yet to effectively address diverse, human-centered tasks. The integration of LLMs with EEG signals to generate personalized diagnoses, treatment plans, and assisted electronic medical records remains an urgent and complex challenge with the following limitations.
@@
Secondly, EEG signals can be viewed as a form of long text data [30], yet the signal sequences from a single channel often exhibit significant redundancy for machine learning approaches [31], [32] [33], [34]. LLMs require a considerable number of tokens [19] to process this redundancy, resulting in unnecessary computational overhead. Additionally, many affective computing tasks rely on EEG data from 32, 64, or more channels, which further diminishes computational efficiency and poses challenges for rapid processing. Therefore, effective data compression is critical to optimizing lightweight LLMs for efficient and responsive deployment.
@@
2 RELATED WORK
The integration of large language models with EEG signals has opened new avenues for affective computing and clinical diagnostics, particularly in emotion recognition, mental health assessment, and neurological disorder diagnosis. This section provides a comprehensive review of current research on LLM-based EEG-assisted diagnostic solutions, summarizes the state of the field, and highlights specific gaps addressed by the proposed EEG Emotion Copilot.
2.1 LLM-Based EEG-Assisted Diagnosis
Recent studies have leveraged LLMs to process EEG signals for diagnostic purposes, capitalizing on their ability to interpret complex data and generate natural language output. In emotion recognition, EEG-GPT [21] demonstrates the potential of LLMs to classify emotional states and motor imagery from raw EEG, though its reliance on large models limits practical deployment. Similarly, NeuroLM [35] integrates EEG and language in a multitask foundation model, improving emotion classification and diagnostic text generation but requiring significant computational resources. Text generation from EEG signals [36] has also been explored, with fine-tuned LLMs producing emotional state descriptions for visual stimuli, yet cross-subject variability remains a challenge.
Beyond emotion, LLMs are applied to mental health and neurological diagnostics. A review by Shang et al. [37] highlights the role of AI in EEG-based diagnosis of epilepsy, Alzheimer’s, and depression, with LLMs generating interpretive reports, although data heterogeneity and constraints on rapid inference still persist. Automated schizophrenia diagnosis using EEG and LLMs for multimodal data fusion shows promise but lacks clinical validation [38]. Specialized LLMs [39] have outperformed neurologists in complex diagnostic tasks, but their EEG integration is limited to textual case processing. Foundation models like LaBraM [40] enable cross-dataset EEG learning for tasks like seizure detection and emotion recognition, yet are not optimized for resource-constrained environments.
Despite these advancements, LLM-based EEG diagnostics face significant hurdles:
  • Computational Cost: Large models (e.g., 14B) require high-end GPUs, hindering clinical deployment.
  • Generalization: EEG variability across subjects, devices, and sessions limits model robustness [15].
  • Privacy: Cloud-based LLMs risk patient data exposure, necessitating local solutions.
  • Clinical Validation: Lack of human-in-the-loop mechanisms and rigorous trials reduces trust in LLM output.
2.2 Lengthy EEG Signal Embedded in Prompt
For specific tasks in LLM prompt engineering, tailored prompts are essential to obtain reasonable answers [41]. Although datasets like Wikitext [26] and Common Crawl [27] contain extensive knowledge, constructing specific datasets is necessary for optimal performance in particular tasks. In most studies on AC, researchers focus on distinguishing emotions using convolutional neural networks (CNNs) or recurrent neural networks (RNNs) [42], with EEG signals primarily used to assess model quality [43], [44]. Recently, LLMs have been required to split EEG sequences into tokens for processing, with computational and memory demands increasing exponentially with input length [45]. This poses significant challenges for LLM deployment, particularly on low-resource devices or in latency-sensitive scenarios. Although data chunking strategies [46] can split long sequences for training and deployment, the EEG data length significantly impacts reasoning effectiveness [47], [48]. Therefore, a reasonable data compression method is necessary to LLMs.
2.3 Model Pruning in Deep Learning
@@
2.4 Assisted Electronic Medical Records
Electronic medical records (EMRs) date back to the 1960s [63]. With advancements in computer science and artificial intelligence (AI), EMRs have evolved from merely storing and managing patient information to incorporating more complex functionalities. Currently AI models are increasingly integrated into diagnosing mental illnesses and detecting internal organ pathologies, enhancing the capabilities of assisted EMRs. The sample of "assisted electronic medical record" is shown in Fig. 2, where the intelligent copilot can process the multimodal diagnostic data and produce accurate results. In this study, we developed a new EEG Emotion Copilot that leverages a lightweight local LLM to perform emotion recognition based on EEG signals, and to further generate the corresponding assisted EMR.
3 METHOD
3.1 Data Structures of Prompt in EEG Emotion Copilot
Fine-tuned LLMs often excel in generating end-to-end outputs, particularly for straightforward classification tasks. For example, determining whether a subject’s emotional state is positive or negative based on EEG data is relatively simple. However, distinguishing more nuanced emotions—such as the nine-class emotional categories in the FACED dataset [29]—or accurately recognizing emotions in real-world scenarios remains a formidable challenge. In datasets like Wikitext and similar corpora, the content primarily consists of declarative sentences (left part of Fig. 3). In contrast, tasks requiring logical reasoning typically store data as question-answer pairs, as illustrated in the middle part of Fig. 3. Despite these advancements, existing data structures face significant limitations when applied to EEG-based emotion diagnosis and treatment planning, particularly in integrating multimodal data and addressing task-specific requirements. To bridge these gaps, we propose a novel data structure that leverages a question-answer system specifically designed for EEG-related tasks, enhancing both interpretability and applicability in clinical and affective computing domains.
@@ 3.1 Data Structures of Prompt in EEG Emotion Copilot
In the data structure of the proposed EEG Emotion Copilot, we initialize the system with the prompt showed in the right part of Fig. 3: "You are an EEG emotion analyzer. I will input the patient’s personal information and EEG signals from specific electrode positions. Please infer the patient’s current emotional state and provide a detailed diagnosis along with personalized treatments based on your knowledge base." Following initialization, we input the subject’s demographic data, including gender, age, and facial features, to assess their emotional state. Next, the acquired EEG signals are processed. As illustrated in Fig. 1, preprocessing steps such as artifact removal and re-referencing are applied to the signals, followed by wavelet transformation for signal compression. These processed signals are then used as input. Specifically, compressed signals are combined with channel information (e.g., Fp1, Cz) into a unified prompt for text input, standardizing the model’s input format. Finally, the structure is completed by adding emotion labels provided by origin dataset and corresponding treatment plans in the second and third sections, respectively.
3.2 Model Pruning in EEG Emotion Copilot
@@
$\mathcal{L}=Num(Qwen2\_layer)=Num(Embedding\_layer)$
$+Num(Qwen2Decoder\_layer)+Num(Lm\_head)=25,$ (1)
where $Num(*)$ denotes the number of respective components: embedding layer, decoder layer, and language modeling head. The input and output of each layer are denoted as $f_{i}^{-}$ and $f_{j}^{+}$, respectively, where $i,j\in\{1,\dots,\mathcal{L}\}$. The superscripts “--” and “$+$” indicate input and output features, respectively.
Next, we construct a layer dependency graph $\mathcal{F}$ according to [64]. This is a $2\mathcal{L}\times 2\mathcal{L}$ symmetric matrix that reflects the structural dependencies between layers:
$&\mathcal{F}\left(f_{i}^{-},f_{j}^{+}\right)=\mathcal{F}\left(f_{j}^{+},f_{i}^{-}\right)\\ &=\begin{cases}1,&\text{if }f_{i}^{-}\leftrightarrow f_{j}^{+}\text{ or }\\ &\quad\left(i=j\text{ and }\operatorname{sch}\left(f_{i}^{-}\right)=\operatorname{sch}\left(f_{j}^{+}\right)\right),\\ 0,&\text{otherwise}.\end{cases}\end{aligned}$ (2)
where $\leftrightarrow$ denotes that two feature maps are directly related and $\text{sch}(p)$ operation is used to determine which parameters in the parameter group $p$ need to be retained. Formula 2 returns a boolean value, where $\text{sch}(f_{i}^{+})=\text{sch}(f_{j}^{-})$ indicates that the two parameter groups are subject to the same pruning scheme.
@@
[…]
$&\mathcal{I}(s_{i})=\sum_{w\in s_{i}}\|w\|_{2}^{2},\\ &S=\{q_{\text{proj}},k_{\text{proj}},v_{\text{proj}},o_{\text{proj}}\}.\end{aligned}\right.$ (4)
$&\mathcal{I}(m_{i})=\sum_{w\in m_{i}}\|w\|_{2}^{2},\\ &M=\{gate_{\text{proj}},up_{\text{proj}},down_{\text{proj}}\}.\end{aligned}\right.$ (5)
[…]
where embed_tokens denotes the input embedding matrix of the Qwen2 model, where each row corresponds to the embedding vector of a token in the vocabulary. This matrix transforms discrete token indices into dense vector representations. $s_{i}$ represents the $i$-th projection matrix within the self-attention module of each Transformer block. Specifically, $s_{i}$ belongs to the set $S=\{q_{\text{proj}},k_{\text{proj}},v_{\text{proj}},o_{\text{proj}}\}$, corresponding to the linear projections of query, key, value, and output in multi-head attention. $m_{i}$ is the $i$-th projection component of the multi-layer perceptron (MLP) block, including the gate, up, and down projections. lm_head refers to the final output linear transformation layer, also known as the language modeling head, which projects the hidden state outputs of the Transformer into vocabulary logits for next-token prediction.
@@
0:  Dependency graph $\mathcal{F}$, parameter group set $\mathcal{P}$
0:  Group set $\mathcal{P}^{\prime}$
1:  Initialize $\mathcal{P}^{\prime}\leftarrow\emptyset$
2:  for $i$in $\mathcal{L}$  do
3:     Create new group $p\leftarrow\{k\}$
4:     Expand group $p$:
5:     while there exists an unvisited node $j$  such that $\mathcal{F}$$(f_{i}^{-},f_{j}^{+})=1$ do
6:        $p\leftarrow p\cup\{j\}$
7:        Mark $j$ as visited
8:     end while
9:     $\mathcal{P}^{\prime}\leftarrow\mathcal{P}^{\prime}\cup p$
10:  end for
11:  return $\mathcal{P}^{\prime}$
Algorithm 2 Pruning Process
0:  Initial model $\mathcal{M}$, Parameter group set $\mathcal{P^{\prime}}$, importance function $\mathcal{I}$, contraction strength $\alpha$
0:  Pruned model $\mathcal{M}^{\prime}$
1:  Calculate group importance:
2:  for $p^{\prime}\in\mathcal{P^{\prime}}$ do
3:     $\mathcal{I}(p^{\prime})\leftarrow\sum_{w\in\mathcal{P^{\prime}}}\|w\|_{2}^{2}$ (See Formula 3, 4, 5, 6)
4:  end for
5:      $\mathcal{P}^{\prime\prime}\leftarrow\text{sort}(\mathcal{P^{\prime}},\text{by }\mathcal{I})$
6:  Select groups for pruning based on importance:
7:  for $p^{\prime\prime}\in\mathcal{P}^{\prime\prime}$ do
8:     if $\mathcal{I}(p^{\prime\prime})$leq $\theta$ is satisfied. ($\theta$ is the threshold)  then
9:        $\mathcal{P^{\prime\prime}}\leftarrow\mathcal{P^{\prime\prime}}\setminus p^{\prime\prime}$
10:     end if
11:  end for
12:  return $\mathcal{M}^{\prime}$
3.3 Model Fine-tuning in EEG Emotion Copilot
@@
To straightforwardly illustrate the fine-tuning process, we take the FACED dataset [29] as a representative example, as shown in Fig. 6. The training procedure is divided into five sequential stages, each consisting of three epochs. In Stages 1 to 4, we employed 250 Hz EEG recordings from 55 subjects in FACED. Interestingly, the model exhibited better performance in Stage 3 than in Stage 4, despite minimal differences in training loss, indicating that extended training may be unnecessary, and potentially harmful—for lightweight architectures. In Stage 5, 1000 Hz EEG data from an additional 68 FACED subjects were incorporated, with the model initialized from the Stage 4 checkpoint. This resulted in richer task-specific responses, further validating the model’s capacity to capture the underlying structure of EEG signals and generalize across varying sampling rates.
3.4 Model Deployment in EEG Emotion Copilot
Given the diverse range of model deployment tools, such as llama.cpp, which enables model quantization based on specific scenarios and conversion into the GGUF format, the model can be seamlessly utilized via Ollama, LLM Studio, Gradio, or other platforms. When deployed with Gradio, this process culminates in the model deployment framework illustrated in Fig. 1.
3.5 Generation of Assisted Electronic Medical Records
Given that the proposed copilot is designed to handle a range of critical tasks—including emotion recognition, diagnosis, and treatment plan formulation—it is imperative to integrate demographic data from diverse cases across various emotional states, alongside their corresponding compressed EEG signals, into the training dataset. In downstream applications, participant data can be structured and fed into the fine-tuned model, enabling the generation of precise diagnostic insights and personalized treatment plans, as illustrated in Fig. 7.